Steam at 100°c is passed into a container of negligible heat capacity containing 20g of ice and 100g of water at 0°c until the ice is completely melted, determine the total mass of water in the container if the specific latent heat of steam is 2300j/g and the S.H.C of water is 4.2j/g/k
1
Expert's answer
2019-11-01T11:40:06-0400
"Q1 = Q2"

Where Q1 - heat of melting ice, Q2- heat of condensation of vapor

Q1= m1 * L1

Q2 = m2 * L2

Where m1 - mass of ice, L1- specific latent heat of melting ice, m2 - mass of water which condensation from vapor(steam), L2 - specific latent heat of steam condensation.

Q1= 20 * 330 = 6600J

Q2= 6600= m2 *2300

m2 = 2.9 g

Summary of water = 100+20+2.9= 122.9 g